Suppose an owner pays $600 million to purchase a hockey team that earns operating profits of $60 million per year. The new owner claims that $400 million of this prices is for the players, which he can depreciate using straight-line depreciation in six years to $100 million. If the team pays corporate profit taxes of 33 percent, how much does the depreciation of the players save the owner each year? Express your answers in millions of dollars. Make sure to round your answers to the nearest 100th decimal points.
Answer #1

Depreciation each year = (P-S)/N = (400m - 100m)/6 = 300m / 6 = 50m

TI with depreciation = Profit - depreciation = 60m - 50 m = 10m

Tax = 0.33 * 10m = 3.3m

TI with out depreciation = 60m

Tax with out depreciation = 0.33 * 60m = 19.80 m

Savings due to depreciation each year = 19.80 m - 3.3m = 16.50m